A threatening note written on 8.5x11 inch paper is pinned to the door of a physics professor. The perpetrator left in a hurry so that when the physics professor finds the paper it is still swinging. Assuming the pin is at the very top center of the paper, what is the period of the paper
Answer #1

I = (1/12)*M*(a^2+b^2) + M*a^2/4 = M*0.0299


d = a/2 =


T = 2*pi*sqrt(I/(M*g*d)

T = 2*pi*sqrt(0.0299/(9.8*0.1397))


T = 0.928 s